### v1

# Learning Multi-Modal Whole-Body Control for Real-World Humanoid Robots

Abstract
We introduce the Masked Humanoid Controller (MHC) for whole-body tracking of target trajectories over arbitrary subsets of humanoid state variables. This enables the realization of whole-body motions from diverse sources such as video, motion capture, and VR, while ensuring balance and robustness against disturbances. The MHC is trained in simulation using a carefully designed curriculum that imitates partially masked motions from a library of behaviors spanning pre-trained policy rollouts, optimized reference trajectories, re-targeted video clips, and human motion capture data. We showcase simulation experiments validating the MHC’s ability to execute a wide variety of behavior from partially-specified target motions. Moreover, we also highlight sim-to-real transfer as demonstrated by real-world trials on the Digit humanoid robot. To our knowledge, this is the first instance of a learned controller that can realize whole-body control of a real-world humanoid for such diverse multi-modal targets.
Keywords: Humanoid Robot Learning, Motion Tracking, Multi-Modal Control
1 Introduction
Humanoid robots hold immense potential as highly capable and adaptive platforms for tackling complex real-world tasks, thanks to their dexterous multi-purpose body structure that mirrors our own. However, the development of versatile and robust whole-body controllers for bipedal humanoids remains a critical challenge in robotics. Traditional approaches involve meticulous manual engineering of separate controllers for different skills such as standing [1], walking  [2], and manipulation  [3], resulting in specialized controllers with limited versatility and adaptability.
A truly versatile whole body humanoid controller should exhibit several key properties. First, it should comprehend and execute target motions specified through multiple input modalities, such as video demonstrations, motion capture data, or high-level locomotion and end-effector targets. Second, the controller should be be robust to dynamic command updates, noisy or inexact inputs, inaccurate simulation parameters, and natural external disturbances. Finally, the controller should be versatile, allowing straightforward extension of its repertoire as new motion examples become available, with no or minimal retraining.
To address these challenges, we train the Masked Humanoid Controller (MHC), a whole-body controller that is able to accommodate target motions specified as future trajectories over full or partial robot poses. This allows the MHC to support the above multi-modality property, for example, following walking trajectories specified solely by desired velocities or imitating arm-only motions extracted from video clips. We train the MHC via reinforcement learning using a carefully designed curriculum and an expansive library of behaviors spanning optimized reference trajectories, re-targeted video clips, and human motion capture data. The tailored curriculum over motion commands and disturbances gradually introduces capabilities with the aim of achieving the above desired robustness and versatility properties.
Figure 1: The Masked Humanoid Controller (MHC) is learned from a dataset of retargeted human motions paired with torso locomotion commands. During training and testing masking can be applied to target motion trajectories to yield masked motion directives that are given to the MHC. The MHC then produces PD setpoints for the whole body in order to track the current motion directive. Training includes domain randomization and force perturbations to facilitate robustness and transfer from simulation to the real robot.
We validate the MHC through simulated benchmarking and showcase qualitative results for sim-to-real transfer using the Digit V3 robot platform. The simulation experiments demonstrate the importance of curriculum and architecture choices and demonstrate generalization to new motions. Our real world experiments demonstrate the robustness of the MHC over a variety of target behaviors such as walking, boxing, and box pickup, while also highlighting current shortcomings in sim-to-real transfer. To our knowledge, this is the first learning-based framework for whole-body real-world humanoid control that can tractably accommodate such a wide spectrum of multi-modal motion directives within a single unified controller.
2 Related Work
Robust Locomotion Control. Recent advancements in robotic locomotion showcase adaptive control strategies that enable robots to navigate complex terrains and perform dynamic maneuvers. Significant progress has been made in the locomotion of quadrupeds, demonstrating enhanced stability [4] [5] and agility [6] in various environments. Fu et al. [7] train a unified whole body manipulation and locomotion controller that attaches an arm on top of a quadraped to perform mobile manipulation tasks. Notable achievements for bipedal locomotion with the Cassie platform (no upper torso or arms) include blind locomotion across multiple gaits and behaviors  [8, 9], locomotion under varying loads  [10], and visually-guided locomotion over irregular terrain using deep learning to process visual input and adapt locomotion strategies accordingly  [11]. More recently a learned controller for blind locomotion on the full-body humanoid Digit was demonstrated across varying terrain and disturbances [12]. These advancements underscore the importance of robust data and simulation technologies in developing and refining control strategies. However, managing transitions between different modes of locomotion, such as standing and walking, remains a significant challenge. Van Marum et al.  [13] addressed this complexity with the development of the SaW controller for Digit, which integrates stability and adaptive walking strategies into a single operational framework that maintains balance against natural disturbances. However, standing and walking alone do not constitute whole-body control. Building on the foundation laid by the SaW controller, our approach enhances it by incorporating comprehensive whole-body control, allowing for more versatile and adaptive humanoid robot behaviors.
Motion Tracking for Simulated Humanoids. Motion capture data has been used extensively in generating motions for simulated characters. Most closely related to our work is the application of reinforcement learning with tracking objective over fully specified target motions [14, 15, 16, 17, 18, 19, 20]. However, tracking rewards alone do not always yield smooth transitions between skills and failure recovery. Recent works augment training with adversarial losses to encourage natural motions during transitions [20, 21] or define an explicit fail state recovery policy [16]. While some works allow for more intuitive under-specified control, they focus on controllers for predefined types of sparsity. Adapting to new sparsity specifications, such as VR, requires retraining [22, 23]. Moreover, discrepancies between simulated and physical environments, such as differences in environmental physics and unrealistic joint definitions in simulated characters, can lead to behaviors that are not physically plausible or efficient when applied to actual robots. Our work aims to both incorporate a wide variety of sparsity types in a single controller and address the simulation to reality gap.
Whole-Body Motion Tracking for Humanoids. Efforts to closely replicate human motion and interaction within robotic frameworks have been significantly enhanced by both simulation techniques and biological inspirations. Early works by Dariush et al.   and recent advancements by Li et al.  [24] have advanced the field of humanoid motion tracking for locomotion, emphasizing the utilization of optimized trajectories and multi-stage training setups to facilitate transfer from simulation to the real world. The latest research by Cheng et al.  [25] employs human motion capture data to achieve partial whole body control, applying imitation constraints on the upper body and relaxing those constraints for lower body. However, all these approaches focus on specific subsets of human motion, such as only upper-body movements or only locomotion control, and may not fully capture the intricate whole-body coordination required for natural and versatile humanoid motion. Our work addresses this limitation by leveraging the MHC framework to enable comprehensive whole-body motion tracking and coordination in humanoids, facilitating more natural and adaptable motion generation.
3 Problem Formulation
Given a dataset of humanoid motions from various behaviors, our objective is to learn a controller which can match target motion directives that are representative of the data distribution. We are particularly interested in supporting partially-specified motion directives e.g. only specifying the upper-body joint trajectories, or just the torso velocity, or both. It is expected that the controller “fills in the blanks” for joints that are not specified by a directive, e.g. details of the lower-body joints. Such a controller can support directives derived from various input modalities. For example, fully-specified directives that cover all humanoid joints can be derived from MoCap data, while joystick commands regarding the root velocity and arm movements correspond to partial directives.
More formally, a motion is a sequence of poses $q_{1:H}$ for a humanoid with $J$ joints over $H$ time steps. Each pose is represented by a tuple $q_{i}=(\theta_{i},{\dot{\theta}}_{i},v_{i},w_{i},b_{i})$, where $\theta_{i}\in\mathbb{R}^{J}$ denotes the joint angles, ${\dot{\theta}}_{i}\in\mathbb{R}^{J}$ denotes the joint angular velocities, $v_{i}\in\mathbb{R}^{2}$ denotes the root planar linear velocities, $w_{i}\in\mathbb{R}$ denotes the turn rate for the humanoid, and $b_{i}\in\mathbb{R}^{3}$ denotes the Euler base orientation in the x-axis, y-axis, and the height of the base from the ground plane. Motion directives are used to specify constraints on a desired motion to be generated. Specifically, a motion directive $d$ is defined as a masked motion sequence represented by $d=(\hat{q}_{1:H},I_{1:H})$, where $\hat{q}_{1:H}$ is a masked motion and $I_{1:H}$ is a sequence of binary masks such that $I_{i}$ indicates which dimensions of pose $q_{i}$ are selected as motion constraints. We follow the convention of setting the masked dimensions of $\hat{q}$ to zero. Note that while the definition of a directive allows for arbitrary sets of state variables to be masked, in practice, we focus training and evaluation masking patterns relevant to the multiple modalities the controller needs to support as described in Section 4.
The Masked Humanoid Controller (MHC) is a controller $\pi$ that at each time step $t$ takes an input containing the current humanoid state $s_{t}=(\theta_{t},\dot{\theta}_{t},\omega_{t})$ and a target motion directive $d$, where $\theta_{t}\in\mathbb{R}^{J}$ and $\dot{\theta}_{t}\in\mathbb{R}^{J}$ are the joint positions and velocities, and $\omega_{t}$ is the orientation in quaternion form. The output of the MHC is an action $a_{t}=\pi(s_{t},d)$, which in our work corresponds to PD setpoints for all robot motors. The objective of the MHC is to select actions such that the future humanoid motion agrees as much as possible with the directive, while maintaining stability and robustness to disturbances. In particular, a major goal of this work is to demonstrate real-world performance of the MHC on the Digit V3 humanoid robot.
4 Masked Humanoid Controller
Due to the lack of low-level supervisory information, we train the MHC via reinforcement learning (RL) in simulation and then transfer the resulting MHC to the real world. Below we describe our choices for the MHC network architecture. Next we describe the training approach, which includes the curriculum of training episodes and domain randomization choices to support sim-to-real transfer. Finally, we describe the reward function used to guide training.
Network Architecture. The MHC model is a Long Short-Term Memory (LSTM) recurrent neural network which takes as input the current humanoid state $s_{t}$ and the next step of the current directive $d_{i}=(\hat{q}_{i},I_{i})$. The MHC first processes the directive with a single-layer feed-forward encoder to produce a 160 dimensional directive embedding. This is concatenated with $s_{t}$ and fed into an LSTM block configured with two 64-unit recurrent layers. A final linear decoding layer outputs an offset for each actuated joint. The MHC action $a_{t}$ is computed by adding this offset to the actuated joint values in $\hat{q}_{i}$, which yields the next PD setpoints. Note that for the masked joints in $d_{i}$, the corresponding values in $\hat{q}_{i}$ are zero and the offset corresponds to the actual PD setpoint. In our work, we use the Digit V3 humanoid, which has 20 actuated joints. The MHC is run at 50Hz to compute PD setpoints which are sent to a PD controller running at 2kHz.
Training Approach. The MHC is trained via the PPO RL algorithm [26] in simulation using the MuJoCo physics engine with a Digit V3 humanoid model. Below we describe the way we generate a motion dataset for creating directives, how we generated training episodes, the curriculum stages used for effective training, and domain randomization to facilitate sim-to-real transfer.
Motion Data Generation. We create a diverse set of reference motions for training the MHC using a combined dataset consisting of human motion capture (MoCap) datasets (AMASS [27], Reallusion [28]), and video demonstrations. To help bridge the human to Digit embodiment gap, we employ an inverse kinematics (IK) based retargeting procedure to map the dataset trajectories to Digit’s kinematic model. For each frame in the dataset, we solve an IK problem for the generalized position vector of digit $q=(\theta,b)$ formulated as a nonlinear program (NLP). We use the IK module in Drake [cite] to set up the costs and constraints associated with the NLP, and SNOPT [cite] as the underlying solver. Critical kinematic feasibility requirements such as locking the stance foot to the ground and respecting the closed kinematic chain topology of Digit’s legs are expressed as constraints in task space. Upper body motion targets that serve more of a stylistic purpose such torso pose, and hand positions are expressed using costs to aid in convergence. Any frames that the IK optimization did not generate a feasible solution for are simply dropped. Linear interpolation is applied to the sequence of states $q_{i}$ and their timestamps $t_{i}$ to produce the final time-parameterized trajectory $q_{1:H}$ used for training.
Episode Generation. Each episode is initialized with the robot in a standing position. Next a random command window length $w$ is generated and a random directive $d$ is drawn from a distribution defined by the curriculum stage. The MHC then cycles through $d$ until reaching $w$ steps. This sampling of a $w$ and $d$ continues until reaching the maximum episode length $e$, which depends on the curriculum stage, or the robot falls. By switching between multiple random directives during an episode, the MHC can learn to smoothly transition between different types of motion. In addition, to encourage robustness, during the execution of each episode we apply perturbations on the torso according to a distribution that depends on the curriculum stage.
Curriculum Stages. We employ a three-stage curriculum that progressively learns locomotion, robustness to perturbations, and whole-body motion tracking. This approach enables the model to gradually acquire complex skills while ensuring stable learning.
(Stage 1: Locomotion) The objective of this stage is to focus learning on developing basic balance and locomotion control, which is a core capability required for more complex behaviors. For this purpose, the episodes involve only randomized locomotion directives, which are partial directives that mask all variables except for those specifying root motion $(v_{i},w_{i},b_{i})$. Instantaneous torso perturbations (80-800N) are also introduced, each affecting a single policy step with a $1\%$ probability. Training lasts for 300 policy steps (6 seconds) with a command window $w\in[40,100]$.
(Stage 2: Stability) Building upon Stage 1, we focus on enhancing the controller’s stability during locomotion. In this stage continuous torso perturbations (20-150N) are applied over windows of 20-50 policy steps. In order to facilitate recovery, we increase episode length to 800 policy steps (16 seconds) and the command window length to $w\in[100,400]$.
(Stage 3: Whole-Body Directives) In this stage, we train the MHC to follow both fully and partially specified directives. Fully-specified directives provide the complete motion reference without any masking, requiring the MHC to perform whole-body imitation. Partially specified directives may take three forms, which correspond to key input modalities: 1) providing only locomotion commands without any upper or lower-body motion specification, corresponding to a joystick input modality, 2) specifying only the upper-body motion while standing in place, upper body imitation from VR or video modalities, or 3) specifying upper-body target motions with locomotion commands, corresponding to a richer VR controller with a locomotion interface. The MHC learns to generate coherent whole-body motions while satisfying the constraints specified by the unmasked components of the directive. Throughout this stage, we maintain the episode length at 800 policy steps and the command window length $w$ within the range [100, 400].
Domain Randomization. To facilitate transfer from simulation to the real robot, we employ dynamics randomization throughout all stages of the curriculum. This involves randomizing various physical parameters of the simulated environment, such as joint damping, link masses, center of mass positions, encoder noise, ground friction, and terrain variations. By exposing the MHC to a wide range of dynamics during training, we aim to learn a policy that can generalize to the dynamics of the real robot without requiring precise system identification or extensive real-world fine-tuning. The randomization ranges are carefully chosen to strike a balance between creating sufficiently diverse training scenarios and maintaining realistic behavior of the simulated robot. Details of the specific randomization ranges used in this work can be found in the the supplementary material.
Reward Design. Our reward function must encourage stability and robustness during locomotion and standing as well as tracking of unmasked joint in the current directive. For locomotion and standing stability we use the reward function used to train a recent state-of-the-art humanoid standing and walking controller [13]. This reward function include three components: 1) Task Reward $r_{task}$ to encourage command following for locomotion directives, 2) Style Reward $r_{style}$ to encourage natural body postures for standing and walking, 3) Regularization Reward $r_{reg}$ to encourage smooth joint motion and low torques. We refer the reader to the original paper for details of these components. To encourage accurate tracking, we include an additional Tracking Reward $r_{track}$. For a given time step let $\theta$ be the current joint values and $\hat{\theta}_{i}$ be the target joint for that step along with the corresponding directive mask $I_{i}$. The corresponding tracking reward of step $t$ is
$r_{track}=\exp\left(-1.5\cdot||(\theta-\hat{\theta}_{i})\cdot I_{i}||\right),$
which decreases with the L2 distance between the target and actual unmasked joints.
The overall reward function depends on whether the directive masks the lower-body joints or not. When the lower-body is masked, e.g. for pure locomotion commands or locomotion combined with upper body tracking, the reward function is $r=r_{task}+r_{style}+r_{reg}+r_{track}$. However, when the lower-body joints are not masked in the directive, the reward function removes the style reward and becomes $r=r_{task}+r_{reg}+r_{track}$. This is done because the motion preferences of the style reward may conflict with the specific lower-body target directives that should be tracked.
5 Experimental Results
In this section, we present our evaluation protocol and simulation experiments to assess the MHC’s performance in executing diverse behaviors from full and partially-specified target motions. We compare the MHC with baseline approaches and analyze the impact of training dataset diversity as well as learning curriculum. Additionally, we discuss qualitative observations and the successful sim-to-real transfer of the MHC through real-world trials on the Digit V3 humanoid robot.
Motion Directive Dataset. We used IK retargetting to generate 75 kinematic motion trajectories selected from our motion sources as described in Section 4. We aimed to select a diverse set of motions, while avoiding aggressive motions that involve highly dynamic behavior such as jumping, kicking, and large sudden swinging of limbs. In order to test models on motions outside of their training sets, we divided the data into three sets, each containing a mix of motions from the original motion sources.
  • setA [20 motions]: Amass Boxing (5), Amass misc (9), Reallusion (2), optimized (2), video (2)
  • setB [20 motions]: Amass Boxing (6), Amass misc (6), Reallusion (5), optimized (3)
  • setC [35 motions]: Amass Boxing (19), Amass misc (12), Reallusion (1), optimized (3)
While each set contains some basic motions related to hand waving and simple upper body movements, the general trend is an increasing level of difficulty as judged by the authors. We note that setC has a number of the most difficult motions, involving highly dynamic motion, such as tennis smashes and difficult boxing moves. The feasibility of these more difficult motions is unclear for our realistic Digit model.
Experiment Setup and Metrics. Each experiment involves evaluating an MHC controller on one of the retargeted motion datasets. For each motion we create 3 motion directives corresponding to a whole body directive, upper body command with lower body standing directive, upper body command with lower body locomotion directive. For each motion directive we evaluate the MHC over 200 episodes, each having a maximum time of 50 seconds (2500 steps). Each episode starts from a randomized point in the directive and then repeatedly cycles through the directive three times before resetting to another randomly selected starting point. Based on these episode runs we evauate the following metrics, which depend on the type of directive.
  • Failure Rate (Fail %): Percentage of episodes that result in failure (i.e. falling) before 50 seconds.
  • Mean per joint positional error ($\mathbb{E}_{\text{MPJPE}}$): Mean positional error for end effectors (hands, elbows, knees, feet) is calculated as the L2 norm of the current Euler position relative to the torso, compared to values from the directive. For partial directives, the error is computed only for unmasked joints. This is done for both partial and fully-specified directives before episode termination.
  • Root Drift (RootΔ): The mean of drift in root position during an episode from its commanded position when following partially or fully-specified directives. Computed as an L2 norm of current root position against expected root position in previous directive step.
Model Fail % $\mathbb{E}_{\text{MPJPE}}$ RootΔ
Locomotion+Offsets 49.213 0.156 0.168
LocomotionTrain 4.800 0.227 0.203
MHC 0.450 0.088 0.134
[…]
Comparison to Baselines. There are no existing approach that handle the masked directive inputs of the MHC. Thus, we have developed two baselines to compare against: 1) Locomotion+Offset. This controller is trained on the first two curriculum stages and only sees the locomotion commands at its input. At inference time, it receives locomotion commands from the directives and for other unmasked joints directly uses the IK offsets in the target motions as the PD setpoints, 2) LocomotionTrain. This is the same as Locomotion+Offset during training and inference, but includes the IK offsets as part of the controller input. The motivation for LocomotionTrain is that by seeing the IK offsets that will be used during motion generation, the controller can adjust to account for the upcoming motion and be more robust.
We train the MHC and two baselines on the SetA data. The evaluation results are provided in Table 1. We see that the MHC outperforms the baselines across all metrics. Most significantly it achieves a failure rate of less than 1% which is significantly lower than the baselines. Interestingly, we see that LocomotionTrain achieves a significantly lower rate than Locomotion+Offsets, showing the utility having the controller learn to prepare for upcoming target motions.
[…]
Table 2: Comparison of Models across Different Datasets. Note that higher failure rates tend to be biased toward smaller (better) values of the other metrics, since those metrics are only averaged over executions before failure.
Figure 2: Comparison of curriculum
stage 1 vs. training from scratch
Figure 3: Comparison of MHC vs. training whole-body directives without stability stage
Overall, these results show that our training approach can generalize to motions outside of the training set. However, more investigation is needed to understand the significance of training and testing with more difficult motions. In particular, the exploration of much longer training runs and continually monitoring of individual motion performance is likely to provide useful insights.
Curriculum Results. To evaluate the effectiveness of our curriculum design, we conducted an ablation study comparing the MHC trained with our complete curriculum against two variants: (1) NoCurriculum, a controller trained without any curriculum, and (2) LocoMask, a controller trained without the stabilization stage. All controllers were trained on setA. The results, shown in Figure 2 and Figure 3, demonstrate the importance of our structured curriculum. The NoCurriculum controller took significantly longer to learn and did not reach the same performance level as the MHC in the initial locomotion stage (Figure 2). The LocoMask controller exhibited greater instability, resulting in shorter average episode lengths (Figure 3). These findings highlight that our staged curriculum, which progressively learns locomotion, stability, and whole-body motion tracking, is very important for achieving our observed robustness and stability.
Sim-to-Real Results. We conduct a real-world demonstration of the MHC using the Digit V3 humanoid robot. Our supplementary video illustrates a number of successful examples and some failures. Some examples of successes include tracking of fully specified boxing directives, partial directives such as robust walking using joystick commands, box pick and place with locomotion + upper body partial directives as well as non-traditional zombie walking gaits. Failure modes included dancing motions that involved balancing on one feet for a substantial amount of time and fully extended arm poses for box pick up motions. Here holding the box outwards, away from the torso caused imbalance leading to a continuous drift from standing posture in order to stabilize. Wheile there is significant room to improvement these sim-to-real results demonstrate that the MHC framework is a promising path toward highly robust and versatile whole-body humanoid control.
Figure 4: A) Locomotion Directives to follow joystick commands; B) Fully-specified directive for a boxing jab displaying whole body torso motion and feet coordination; C) Masked partial directives for upper body motion tracking and lower body locomotion commands to achieve the task of moving to a location, picking up the box, moving while holding the box and placing it at a location.
6 Limitations
One of the major current limitations is a significant sim-to-real gap. This is particularly prevalent for motions where the feet are placed widely or where one foot remains in the air for an extended period. Our hypothesis is that this sim-to-real limitation requires research that is more generally focused on sim-to-real transfer, possibly via the incorporation of real-world data. The MHC has been a useful vehicle for highlighting these gaps. Another limitation of our current training approach is that we only consider a limited set of masking patterns during training. Supporting fully general masking may lead to additional robustness and open up new possibilities for authoring behaviors. Finally, the MHC is currently does not account for interaction with external objects, such as the box, during a box pickup. Explicitly incorporating objects to be manipulated into the MHC input and training is an important direction for supporting practical applications.
Appendix A Dynamics Randomization
All models in our work were trained with a fixed dynamics randomization setup in order to transfer from simulation to real without learning the exact dynamics of the system. The ranges for all elements are given in Table 3.
Element Range(%)
Joint Damping [-50, 250]
Body Mass [-25, 25]
CoM Position [-5, 5]
Spring Stiffness [-10, 10]
Friction [-20, 20]
Range(radians)
Encoder Noise [-0.05, 0.05]
Ground Slope [-0.05, 0.05]
Table 3: Dynamics Randomization Ranges.
Appendix B Implementation Details
MHC has been trained using actor critic based PPO algorithm. Table 4 specifies the hyperparameters used for each training cycle independent of the curriculum.
Hyperparameter Value
Learning Rate (Actor & Critic) 3e-4
Buffer Size 50000
Number of Epochs 5
Batch Size (of episodes) 8
PPO Log-Clip Range 0.2
Maximum Gradient Norm 0.05
Discount Factor (Gamma) 0.95
GAE Lambda 0.95
Table 4: PPO Hyperparameters
Appendix C Reward structure
Table 5 gives a detailed view of our complete reward function. Note that style rewards are only ignored when following a fully-specified directive denoted by $I_{i}==1$.
Following notation has been used:
  • $i$ indicates current time step
  • $I_{i}==1$ indicated if all joints are being mimicked and its a whole body motion directive;
  • $v_{xy}$ = base linear velocity;
  • $s$ = current robot parameter;
  • $\boldsymbol{b}$ = base acceleration;
  • $c_{feet}$ = default foot position xyz and orientation rpy;
  • $qd(\cdot)$ = quaternion distance function;
  • $n_{c}$ = number of feet in contact with ground;
  • $\mathds{1}_{td}$ = boolean variable indicating a foot touchdown in the current timestep;
Reward Term Definition Weight Active if $I_{i}==1$
if $v_{i}==0$ [Standing] else [Walking]
Task $x,y$ velocity $\exp\left(-5\cdot(v_{xy})\right)$ $\exp\left(-5\cdot(v_{xy}-v_{i})^{2}\right)$ 0.15
Tracking Reward $\exp\left(-1.5\cdot||(\theta-\hat{\theta}_{i})\cdot I_{i}||\right)$ 0.2
Task, Style Yaw orient $\exp\left(-300\cdot qd(\omega_{z},w_{i})\right)$ 0.1
Roll, pitch orient $\exp\left(-30\cdot qd\big{(}\omega_{xy},(b_{i}[xy])\big{)}\right)$ 0.2
Base height $\exp\left(-20\cdot\big{|}s_{z}-(b_{i}[z])\big{|}\right)$ 0.05
Style Feet orientation $\exp\left(-\sum|s_{\text{feet, rpy}}-c_{\text{feet, rpy}}|\right)$ 0.05
Feet position $\exp\left(-3\cdot|S_{\text{feet, xyz}}-c_{\text{feet, xyz}}|\right)$ 1 0.05
Feet airtime 1 $\sum_{f\in(l,r)}(t_{air,f}-0.4)*\mathds{1}_{td,f}$ 1.0
Feet contact 1 $\begin{cases}1&\text{if }n_{c,t^{*}}=1\text{ for any }t^{*}\in[t-0.2,t]\\ 0&\text{else}\end{cases}$ 0.1
Arm $\exp\left(-3\cdot\big{|}\big{|}(\theta[\hat{J}]\text{ where }\hat{J}\in\text{ arms})\big{|}\big{|}\right)$ 0.03
Reg Base acceleration $\exp\left(-0.01\cdot\sum|b_{xyz}|\right)$ 0.1
Action difference $\exp\left(-0.02\cdot\sum|a_{t}-a_{t-1}|\right)$ 0.02
Torque $\exp\left(-0.02\cdot\frac{1}{N}\sum|t_{\text{motor}}|/t_{\text{max}}\right)$ 0.02
Table 5: Reward Terms

### v4

@@
 Abstract
-We introduce the Masked Humanoid Controller (MHC) for whole-body tracking of target trajectories over arbitrary subsets of humanoid state variables.
-This enables the realization of whole-body motions from diverse sources such as video, motion capture, and VR, while ensuring balance and robustness against disturbances.
-The MHC is trained in simulation using a carefully designed curriculum that imitates partially masked motions from a library of behaviors spanning pre-trained policy rollouts, optimized reference trajectories, re-targeted video clips, and human motion capture data.
-We showcase simulation experiments validating the MHC’s ability to execute a wide variety of behavior from partially-specified target motions.
-Moreover, we also highlight sim-to-real transfer as demonstrated by real-world trials on the Digit humanoid robot.
-To our knowledge, this is the first instance of a learned controller that can realize whole-body control of a real-world humanoid for such diverse multi-modal targets.
-Keywords: Humanoid Robot Learning, Motion Tracking, Multi-Modal Control
-1 Introduction
+A major challenge in humanoid robotics is designing a unified interface for commanding diverse whole-body behaviors, from precise footstep sequences to partial-body mimicry and joystick teleoperation.
+We introduce the Masked Humanoid Controller (MHC), a learned whole-body controller that exposes a simple yet expressive interface: the specification of masked target trajectories over selected subsets of the robot’s state variables.
+This unified abstraction allows high-level systems to issue commands in a flexible format that accommodates multi-modal inputs such as optimized trajectories, motion capture clips, re-targeted video, and real-time joystick signals.
+The MHC is trained in simulation using a curriculum that spans this full range of modalities, enabling robust execution of partially specified behaviors while maintaining balance and disturbance rejection.
+We demonstrate the MHC both in simulation and on the real-world Digit V3 humanoid, showing that a single learned controller is capable of executing such diverse whole-body commands in the real world through a common representational interface.
+Project Webpagehttps://masked-humanoid.github.io/mhc/
+Introduction
 Humanoid robots hold immense potential as highly capable and adaptive platforms for tackling complex real-world tasks, thanks to their dexterous multi-purpose body structure that mirrors our own.
-However, the development of versatile and robust whole-body controllers for bipedal humanoids remains a critical challenge in robotics.
-Traditional approaches involve meticulous manual engineering of separate controllers for different skills such as standing [1], walking  [2], and manipulation  [3], resulting in specialized controllers with limited versatility and adaptability.
-A truly versatile whole body humanoid controller should exhibit several key properties.
-First, it should comprehend and execute target motions specified through multiple input modalities, such as video demonstrations, motion capture data, or high-level locomotion and end-effector targets.
-Second, the controller should be be robust to dynamic command updates, noisy or inexact inputs, inaccurate simulation parameters, and natural external disturbances.
-Finally, the controller should be versatile, allowing straightforward extension of its repertoire as new motion examples become available, with no or minimal retraining.
-To address these challenges, we train the Masked Humanoid Controller (MHC), a whole-body controller that is able to accommodate target motions specified as future trajectories over full or partial robot poses.
-This allows the MHC to support the above multi-modality property, for example, following walking trajectories specified solely by desired velocities or imitating arm-only motions extracted from video clips.
+However, development of versatile and robust whole-body controllers for bipedal humanoids remains a critical challenge in robotics.
+Traditional approaches involve meticulous manual engineering of separate controllers for different skills such as standing (Crowley et al.
+2023), walking (Castillo et al.
+2021), manipulation (Dao et al.
+2023) and mimicry (He et al.
+2024a), resulting in specialized controllers with limited versatility and adaptability.
+While effective in isolation, these controllers lack versatility, are difficult to extend, and rarely support seamless reuse across tasks.
+A truly versatile whole-body humanoid controller should exhibit several key properties.
+First, it should be able to robustly stand, walk, and mimic whole and partial-body motions.
+Second, it should comprehend and execute these capabilities through multiple input modalities, including combinations of high-level velocity commands, end-effector targets, video demonstrations, or motion capture data.
+Third, the controller should be robust to dynamic command updates, noisy or inexact inputs, inaccurate simulation parameters, and external disturbances.
+Finally, the controller should allow for straightforward extension of its repertoire as new motion examples become available, with no or minimal retraining.
+A key challenge is the lack of a unified representation for specifying robot behaviors.
+Tasks may be defined through velocity commands, end-effector targets, human demonstrations, or partial-body trajectories, yet most systems rely on distinct processing pipelines for each modality.
+This fragmentation slows application development, restricts interoperability, and complicates the transfer of policies from simulation to real hardware.
+To address these challenges, we train the Masked Humanoid Controller (MHC), a whole-body controller that provides a unified representational interface for multi-modal humanoid control.
+MHC represents target behaviors as future pose trajectories with optional masking, allowing the same framework to accommodate velocity-based walking, arm-only imitation, whole body imitation or combined loco-manipulation tasks.
 We train the MHC via reinforcement learning using a carefully designed curriculum and an expansive library of behaviors spanning optimized reference trajectories, re-targeted video clips, and human motion capture data.
 The tailored curriculum over motion commands and disturbances gradually introduces capabilities with the aim of achieving the above desired robustness and versatility properties.
-Figure 1: The Masked Humanoid Controller (MHC) is learned from a dataset of retargeted human motions paired with torso locomotion commands.
-During training and testing masking can be applied to target motion trajectories to yield masked motion directives that are given to the MHC.
+Figure 1: The Masked Humanoid Controller (MHC) is learned from a dataset of re-targeted human motions paired with torso locomotion commands, including standing.
+During training and testing, masking can be applied to target motion trajectories to yield masked motion directives that are given to the MHC.
 The MHC then produces PD setpoints for the whole body in order to track the current motion directive.
 Training includes domain randomization and force perturbations to facilitate robustness and transfer from simulation to the real robot.
-We validate the MHC through simulated benchmarking and showcase qualitative results for sim-to-real transfer using the Digit V3 robot platform.
-The simulation experiments demonstrate the importance of curriculum and architecture choices and demonstrate generalization to new motions.
-Our real world experiments demonstrate the robustness of the MHC over a variety of target behaviors such as walking, boxing, and box pickup, while also highlighting current shortcomings in sim-to-real transfer.
-To our knowledge, this is the first learning-based framework for whole-body real-world humanoid control that can tractably accommodate such a wide spectrum of multi-modal motion directives within a single unified controller.
-2 Related Work
+We validate the MHC through simulated benchmarking and showcase qualitative and quantitative results for sim-to-real transfer using the Digit V3 robot platform.
+The simulation experiments demonstrate the importance of the curriculum, architecture choices and generalization to new motions.
+Our real-world experiments demonstrate the robustness of the MHC across a variety of target behaviors such as walking, boxing, and box loco-manipulation, while also highlighting current shortcomings in sim-to-real transfer.
+Related Work
+Motion Tracking for Simulated Humanoids.
+Motion capture data have been used extensively for generating motions of simulated characters.
+Most closely related to our work is the application of reinforcement learning with tracking rewards over fully specified target motions (Won et al.
+2022; Luo et al.
+2023; Merel et al.
+2018; Dou et al.
+2023; Wagener et al.
+2022; Peng et al.
+2022).
+However, tracking rewards alone do not always yield smooth transitions between skills and failure recovery.
+Recent works augment training with adversarial losses to encourage natural motions during transitions (Peng et al.
+2022; Tessler et al.
+2023) or define an explicit fail state recovery policy (Luo et al.
+2023).
+While some approaches offer more intuitive control for under-specified tasks, they often focus on predefined types of input trajectory sparsity, and adapting to new sparsity specifications, such as VR, require retraining (Cerón et al.
+2023; Du et al.
+2023).
+Notably, existing models designed for physically realistic animation use the SMPL skeleton in IsaacSIM, such as PHC (Luo et al.
+2023), and ASE (Peng et al.
+2022), and require substantial engineering effort and fine-tuning to adapt to a real humanoid in the MuJoCo setup with dynamics randomization.
+Moreover, discrepancies between simulated and physical environments—such as differences in environmental physics and unrealistic joint definitions in simulated characters—can result in behaviors that are not physically plausible or efficient for actual robots.
+Work such as Versatile Motion Priors (VMP) (Serifi et al.
+2024) learns latent motion embeddings from unstructured human motion and trains dynamics-aware policies that can generalize to diverse behaviors on real humanoids, showing promise for bridging the simulation-to-reality gap.
+Recent work on the Masked Humanoid Controller (MHC) (Shrestha et al.
+2024) addresses these challenges by learning a unified low-level policy from sparsely masked motion data, enabling robust tracking, smooth transitions, and generalization across diverse, under-specified input modalities.
+Our work similarly aims to incorporate a wide variety of sparsity types in a single controller while focusing more on addressing the simulation to reality gap.
 Robust Locomotion Control.
-Recent advancements in robotic locomotion showcase adaptive control strategies that enable robots to navigate complex terrains and perform dynamic maneuvers.
-Significant progress has been made in the locomotion of quadrupeds, demonstrating enhanced stability [4] [5] and agility [6] in various environments.
-Fu et al.
-[7] train a unified whole body manipulation and locomotion controller that attaches an arm on top of a quadraped to perform mobile manipulation tasks.
-Notable achievements for bipedal locomotion with the Cassie platform (no upper torso or arms) include blind locomotion across multiple gaits and behaviors  [8, 9], locomotion under varying loads  [10], and visually-guided locomotion over irregular terrain using deep learning to process visual input and adapt locomotion strategies accordingly  [11].
-More recently a learned controller for blind locomotion on the full-body humanoid Digit was demonstrated across varying terrain and disturbances [12].
+Standing, walking, and transitioning from one to the other present distinct challenges in stability and adaptability, driving recent advances in adaptive control strategies that enable robots to navigate complex terrains and perform dynamic maneuvers.
+Significant progress has been made in the locomotion of quadrupeds, demonstrating enhanced stability (Lee et al.
+2020; Da et al.
+2021) and agility (Jenelten et al.
+2019) in various environments.
+(Fu et al.
+2023) train a unified whole-body manipulation and locomotion controller that attaches an arm on top of a quadruped to perform mobile manipulation tasks.
+Notable achievements for bipedal locomotion with the Cassie platform (no upper torso or arms) include blind locomotion across multiple gaits and behaviors (Siekmann et al.
+2021; Li et al.
+2021), locomotion under varying loads (Dao et al.
+2022), and visually-guided locomotion over irregular terrain using deep learning to process visual input and adapt locomotion strategies accordingly (Duan et al.
+2023).
+More recently a learned controller for blind locomotion on the full-body humanoid Digit was demonstrated across varying terrain and disturbances (Radosavovic et al.
+2023).
 These advancements underscore the importance of robust data and simulation technologies in developing and refining control strategies.
-However, managing transitions between different modes of locomotion, such as standing and walking, remains a significant challenge.
+However, managing transitions between different modes of locomotion, such as standing and walking, remained a significant challenge.
 Van Marum et al.
-[13] addressed this complexity with the development of the SaW controller for Digit, which integrates stability and adaptive walking strategies into a single operational framework that maintains balance against natural disturbances.
+(van Marum et al.
+2024) addressed this complexity with the development of the SaW controller for Digit, which integrates stability and adaptive walking strategies into a single operational framework that maintains balance against natural disturbances.
 However, standing and walking alone do not constitute whole-body control.
-Building on the foundation laid by the SaW controller, our approach enhances it by incorporating comprehensive whole-body control, allowing for more versatile and adaptive humanoid robot behaviors.
-Motion Tracking for Simulated Humanoids.
-Motion capture data has been used extensively in generating motions for simulated characters.
-Most closely related to our work is the application of reinforcement learning with tracking objective over fully specified target motions [14, 15, 16, 17, 18, 19, 20].
-However, tracking rewards alone do not always yield smooth transitions between skills and failure recovery.
-Recent works augment training with adversarial losses to encourage natural motions during transitions [20, 21] or define an explicit fail state recovery policy [16].
-While some works allow for more intuitive under-specified control, they focus on controllers for predefined types of sparsity.
-Adapting to new sparsity specifications, such as VR, requires retraining [22, 23].
-Moreover, discrepancies between simulated and physical environments, such as differences in environmental physics and unrealistic joint definitions in simulated characters, can lead to behaviors that are not physically plausible or efficient when applied to actual robots.
-Our work aims to both incorporate a wide variety of sparsity types in a single controller and address the simulation to reality gap.
-Whole-Body Motion Tracking for Humanoids.
+Mimicry for Humanoids.
 Efforts to closely replicate human motion and interaction within robotic frameworks have been significantly enhanced by both simulation techniques and biological inspirations.
 Early works by Dariush et al.
-and recent advancements by Li et al.
-[24] have advanced the field of humanoid motion tracking for locomotion, emphasizing the utilization of optimized trajectories and multi-stage training setups to facilitate transfer from simulation to the real world.
-The latest research by Cheng et al.
-[25] employs human motion capture data to achieve partial whole body control, applying imitation constraints on the upper body and relaxing those constraints for lower body.
-However, all these approaches focus on specific subsets of human motion, such as only upper-body movements or only locomotion control, and may not fully capture the intricate whole-body coordination required for natural and versatile humanoid motion.
-Our work addresses this limitation by leveraging the MHC framework to enable comprehensive whole-body motion tracking and coordination in humanoids, facilitating more natural and adaptable motion generation.
-3 Problem Formulation
+(Dariush et al.
+2008) and recent advancements by Li et al.
+(Li et al.
+2024) have advanced the field of humanoid motion tracking for locomotion, emphasizing the utilization of optimized trajectories and multi-stage training setups to facilitate transfer from simulation to the real world.
+In concurrent work, He et al.
+(He et al.
+2024a) have developed a controller trained on the AMASS dataset (Mahmood et al.
+2019) to learn through mimicry, another key modality of control.
+While this approach effectively performs whole-body control, it only covers the range of locomotion capabilities present in the dataset.
+As a result, the system’s performance is constrained by the locomotion patterns and upper body limitations observed in the mimicry data.
+They do not demonstrate robust joystick-style locomotion capabilities that are required for loco-manipulation tasks.
+Upper-Body Motion Tracking.
+Recent research by Cheng et al.
+(Cheng et al.
+2024) covers the modality of partial-body control by performing upper-body mimicry from target motions and joystick-based lower-body control.
+It cannot however handle the modalities of standing or whole-body mimicry.
+It performs a continuous step-in-place motion when commanded for 0 velocity and lacks height control beyond the region of its nominal height.
+These are properties that are crucial for whole-body mimicry when considering lower body.
+It also necessitates the need for commanding a motion for upper-body even while performing simple locomotion tasks.
+An unconstrained locomotion controller would be able to learn optimal arm swing motions that are speed-dependent and perform better under disturbances as the arms would have more mobility when left unconstrained.
+The vast majority of the above approaches focus on a single input modality, such as walking, standing, partial-body mimicry, or whole-body mimicry.
+In order to jointly handle all these input modalities, these approaches would need to either switch between controllers or have an interpreter that can generate a whole-body motion for different speeds and optimal arm positions for the commanded speed on the fly.
+This can result in significant added complexity.
+Concurrent work such as HOVER (He et al.
+2024b) demonstrates a unified policy architecture capable of handling multiple control modes without explicit switching by distilling diverse skills into a single representation space.
+Our MHC follows a similar philosophy, integrating locomotion and comprehensive whole-body motion tracking in humanoids into a single unified framework.
+It enables natural and adaptable motion generation that a user can control based on the scenario at hand with no overhead of policy switching or complex interpreters.
+Problem Formulation
 Given a dataset of humanoid motions from various behaviors, our objective is to learn a controller which can match target motion directives that are representative of the data distribution.
@@
 Such a controller can support directives derived from various input modalities.
-For example, fully-specified directives that cover all humanoid joints can be derived from MoCap data, while joystick commands regarding the root velocity and arm movements correspond to partial directives.
+For example, fully-specified directives specifying all joints can be derived from MoCap data, while joystick commands regarding the root velocity and arm movements correspond to partial directives.
 More formally, a motion is a sequence of poses $q_{1:H}$ for a humanoid with $J$ joints over $H$ time steps.
@@
 We follow the convention of setting the masked dimensions of $\hat{q}$ to zero.
-Note that while the definition of a directive allows for arbitrary sets of state variables to be masked, in practice, we focus training and evaluation masking patterns relevant to the multiple modalities the controller needs to support as described in Section 4.
-The Masked Humanoid Controller (MHC) is a controller $\pi$ that at each time step $t$ takes an input containing the current humanoid state $s_{t}=(\theta_{t},\dot{\theta}_{t},\omega_{t})$ and a target motion directive $d$, where $\theta_{t}\in\mathbb{R}^{J}$ and $\dot{\theta}_{t}\in\mathbb{R}^{J}$ are the joint positions and velocities, and $\omega_{t}$ is the orientation in quaternion form.
+Note that while the definition of a directive allows for arbitrary sets of state variables to be masked, in practice, we focus training and evaluation on masking patterns relevant to the multiple input modalities the controller needs to support as described in Section Masked Humanoid Controller.
+The Masked Humanoid Controller (MHC) is a controller $\pi$ that at each time step $t$ takes an input containing the current humanoid proprioceptive state $s_{t}=(r\theta_{t},\dot{r\theta}_{t},\omega_{t})$ and a target motion directive $d=(\hat{q}_{1:H},I_{1:H})$, where $r\theta_{t}\in\mathbb{R}^{J}$ and $\dot{r\theta}_{t}\in\mathbb{R}^{J}$ are the joint positions and velocities, and $\omega_{t}$ is the torso orientation in quaternion form.
+Note that $s_{t}$ is composed of the actual robot states while $d$ is composed of the motion being commanded.
 The output of the MHC is an action $a_{t}=\pi(s_{t},d)$, which in our work corresponds to PD setpoints for all robot motors.
-The objective of the MHC is to select actions such that the future humanoid motion agrees as much as possible with the directive, while maintaining stability and robustness to disturbances.
-In particular, a major goal of this work is to demonstrate real-world performance of the MHC on the Digit V3 humanoid robot.
-4 Masked Humanoid Controller
-Due to the lack of low-level supervisory information, we train the MHC via reinforcement learning (RL) in simulation and then transfer the resulting MHC to the real world.
-Below we describe our choices for the MHC network architecture.
-Next we describe the training approach, which includes the curriculum of training episodes and domain randomization choices to support sim-to-real transfer.
-Finally, we describe the reward function used to guide training.
-Network Architecture.
-The MHC model is a Long Short-Term Memory (LSTM) recurrent neural network which takes as input the current humanoid state $s_{t}$ and the next step of the current directive $d_{i}=(\hat{q}_{i},I_{i})$.
+The objective of the MHC is to select actions which aligns the future humanoid motion with the commanded directive, while maintaining stability and robustness to disturbances.
+This formulation allows a single controller to generalize across diverse input modalities, supporting both robotic research workflows and artistic motion design.
+For example, specifying only torso velocity and arm trajectories enables the controller to autonomously generate stable leg motions, while fully-specified directives from motion capture ensure high-fidelity imitation.
+This unified representation removes the need for multiple specialized controllers or interpreters.
+Masked Humanoid Controller
+Due to the lack of low-level supervisory information, we train the MHC via reinforcement learning (RL) in simulation before transferring it to the real world.
+The MHC is trained using the PPO algorithm in a MuJoCo physics engine simulation with a Digit V3 humanoid model.
+Below we detail our training approach which involves several key components: motion dataset generation, MHC network architecture, a curriculum of training episodes, domain randomization for sim-to-real transfer, and our mask aligned reward function.
+Motion Data Generation
+We create a diverse set of reference motions for training the MHC using a combined dataset consisting of human motion capture (MoCap) datasets (AMASS (Mahmood et al.
+2019), Reallusion (Reallusion 2022)), video demonstrations, and optimized trajectories designed using spring mass model of Digit.
+To help bridge the human to Digit embodiment gap, we employ an Inverse Kinematics (IK) based retargeting procedure to map the dataset trajectories to Digit’s kinematic model.
+For each frame in the dataset, we solve an IK problem for the generalized position vector of digit $q=(\theta,b)$ formulated as a nonlinear program (NLP).
+We use the IK module in Drake (Tedrake and the Drake Development Team 2019) to set up the costs and constraints associated with the NLP, and SNOPT (Gill et al.
+2005) as the underlying solver.
+Critical kinematic feasibility requirements such as locking the stance foot to the ground and respecting the closed kinematic chain topology of Digit’s legs are expressed as constraints in task space.
+Upper body motion targets that serve more of a stylistic purpose such as torso pose, and hand positions are expressed using costs to aid in convergence.
+Any frames that the IK optimization did not generate a feasible solution for are simply dropped.
+Linear interpolation is applied to the sequence of states $q_{i}$ and their timestamps $t_{i}$ to produce the final time-parameterized trajectory $q_{1:H}$ used for training.
+We note that there are some discrepancies in retargeted motions when MoCap trajectories are not kinematically feasible for the Digit robot.
+This is due to Digit’s one-dimensional rotational joints (limited to $360^{\circ}$ rotation) vs MoCap’s 3D spherical joint models.
+The retargeting process uses inverse kinematics (IK) with cost-based optimization to find the best approximate solution within these constraints.
+Network Architecture
+The MHC model is a Long Short-Term Memory (LSTM) (Hochreiter and Schmidhuber 1997) recurrent neural network which takes as input the current humanoid state $s_{t}$ and the next step of the current directive $d_{i}=(\hat{q}_{i},I_{i})$.
 The MHC first processes the directive with a single-layer feed-forward encoder to produce a 160 dimensional directive embedding.
-This is concatenated with $s_{t}$ and fed into an LSTM block configured with two 64-unit recurrent layers.
+This is concatenated with $s_{t}$ and fed into an LSTM block configured with two 64-unit recurrent layers to capture temporal information.
 A final linear decoding layer outputs an offset for each actuated joint.
@@
 The MHC is run at 50Hz to compute PD setpoints which are sent to a PD controller running at 2kHz.
-Training Approach.
-The MHC is trained via the PPO RL algorithm [26] in simulation using the MuJoCo physics engine with a Digit V3 humanoid model.
-Below we describe the way we generate a motion dataset for creating directives, how we generated training episodes, the curriculum stages used for effective training, and domain randomization to facilitate sim-to-real transfer.
-Motion Data Generation.
-We create a diverse set of reference motions for training the MHC using a combined dataset consisting of human motion capture (MoCap) datasets (AMASS [27], Reallusion [28]), and video demonstrations.
-To help bridge the human to Digit embodiment gap, we employ an inverse kinematics (IK) based retargeting procedure to map the dataset trajectories to Digit’s kinematic model.
-For each frame in the dataset, we solve an IK problem for the generalized position vector of digit $q=(\theta,b)$ formulated as a nonlinear program (NLP).
-We use the IK module in Drake [cite] to set up the costs and constraints associated with the NLP, and SNOPT [cite] as the underlying solver.
-Critical kinematic feasibility requirements such as locking the stance foot to the ground and respecting the closed kinematic chain topology of Digit’s legs are expressed as constraints in task space.
-Upper body motion targets that serve more of a stylistic purpose such torso pose, and hand positions are expressed using costs to aid in convergence.
-Any frames that the IK optimization did not generate a feasible solution for are simply dropped.
-Linear interpolation is applied to the sequence of states $q_{i}$ and their timestamps $t_{i}$ to produce the final time-parameterized trajectory $q_{1:H}$ used for training.
-Episode Generation.
+Training Episode Generation
 Each episode is initialized with the robot in a standing position.
-Next a random command window length $w$ is generated and a random directive $d$ is drawn from a distribution defined by the curriculum stage.
+Next a random command window length $w$ is generated and a random directive $d$ is drawn from a distribution defined by the curriculum stage (see below).
 The MHC then cycles through $d$ until reaching $w$ steps.
-This sampling of a $w$ and $d$ continues until reaching the maximum episode length $e$, which depends on the curriculum stage, or the robot falls.
-By switching between multiple random directives during an episode, the MHC can learn to smoothly transition between different types of motion.
+This sampling of a $w$ and $d$ continues until the robot falls or reaches the maximum episode length $e$, which depends on the curriculum stage.
+By switching between multiple random directives during an episode, the MHC can learn to smoothly transition between different types of motions.
 In addition, to encourage robustness, during the execution of each episode we apply perturbations on the torso according to a distribution that depends on the curriculum stage.
-Curriculum Stages.
-We employ a three-stage curriculum that progressively learns locomotion, robustness to perturbations, and whole-body motion tracking.
-This approach enables the model to gradually acquire complex skills while ensuring stable learning.
-(Stage 1: Locomotion) The objective of this stage is to focus learning on developing basic balance and locomotion control, which is a core capability required for more complex behaviors.
+Curriculum Stages
+We employ a three-stage curriculum that progressively focuses training on locomotion, robustness to perturbations, and whole-body motion tracking.
+This enables gradual acquisition of complex skills while ensuring stable learning.
+Stage 1 - Locomotion: The objective of this stage is to focus learning on developing basic balance and locomotion control, which is a core capability required for more complex behaviors.
 For this purpose, the episodes involve only randomized locomotion directives, which are partial directives that mask all variables except for those specifying root motion $(v_{i},w_{i},b_{i})$.
@@
 Training lasts for 300 policy steps (6 seconds) with a command window $w\in[40,100]$.
-(Stage 2: Stability) Building upon Stage 1, we focus on enhancing the controller’s stability during locomotion.
+Stage 2 - Stability: Building upon Stage 1, we focus on enhancing the controller’s stability during locomotion.
 In this stage continuous torso perturbations (20-150N) are applied over windows of 20-50 policy steps.
 In order to facilitate recovery, we increase episode length to 800 policy steps (16 seconds) and the command window length to $w\in[100,400]$.
-(Stage 3: Whole-Body Directives) In this stage, we train the MHC to follow both fully and partially specified directives.
+Stage 3 - Whole-Body Directives: In this stage, we train the MHC to follow both fully and partially specified directives.
 Fully-specified directives provide the complete motion reference without any masking, requiring the MHC to perform whole-body imitation.
@@
 Throughout this stage, we maintain the episode length at 800 policy steps and the command window length $w$ within the range [100, 400].
-Domain Randomization.
+Domain Randomization
 To facilitate transfer from simulation to the real robot, we employ dynamics randomization throughout all stages of the curriculum.
-This involves randomizing various physical parameters of the simulated environment, such as joint damping, link masses, center of mass positions, encoder noise, ground friction, and terrain variations.
+This involves randomizing various physical parameters of the simulated environment, such as joint damping, link masses, center of mass positions, encoder noise, ground friction.
 By exposing the MHC to a wide range of dynamics during training, we aim to learn a policy that can generalize to the dynamics of the real robot without requiring precise system identification or extensive real-world fine-tuning.
 The randomization ranges are carefully chosen to strike a balance between creating sufficiently diverse training scenarios and maintaining realistic behavior of the simulated robot.
-Details of the specific randomization ranges used in this work can be found in the the supplementary material.
-Reward Design.
-Our reward function must encourage stability and robustness during locomotion and standing as well as tracking of unmasked joint in the current directive.
-For locomotion and standing stability we use the reward function used to train a recent state-of-the-art humanoid standing and walking controller [13].
-This reward function include three components: 1) Task Reward $r_{task}$ to encourage command following for locomotion directives, 2) Style Reward $r_{style}$ to encourage natural body postures for standing and walking, 3) Regularization Reward $r_{reg}$ to encourage smooth joint motion and low torques.
+We use the same randomization ranges as prior work on Digit locomotion training (van Marum et al.
+2024).
+Reward Design
+Our reward function must encourage stability and robustness during locomotion and standing as well as tracking of unmasked joints in the current directive.
+For locomotion and standing stability we build upon the reward function used to train a recent state-of-the-art humanoid standing and walking controller (van Marum et al.
+2024).
+That reward function includes three components: 1) Task Reward $r_{task}$ to encourage command following for locomotion directives, 2) Style Reward $r_{style}$ to encourage natural body postures for standing and walking, 3) Regularization Reward $r_{reg}$ to encourage smooth joint motion and low torques.
 We refer the reader to the original paper for details of these components.
-To encourage accurate tracking, we include an additional Tracking Reward $r_{track}$.
-For a given time step let $\theta$ be the current joint values and $\hat{\theta}_{i}$ be the target joint for that step along with the corresponding directive mask $I_{i}$.
-The corresponding tracking reward of step $t$ is
-$r_{track}=\exp\left(-1.5\cdot||(\theta-\hat{\theta}_{i})\cdot I_{i}||\right),$
-which decreases with the L2 distance between the target and actual unmasked joints.
+We build on these components and add a Tracking Reward $r_{track}$.
+[…]
+where $\theta$ is the current joint values and $\hat{\theta}_{i}$ is the target joint for that step along with the corresponding directive mask $I_{i}$.
 The overall reward function depends on whether the directive masks the lower-body joints or not.
@@
 This is done because the motion preferences of the style reward may conflict with the specific lower-body target directives that should be tracked.
-5 Experimental Results
-In this section, we present our evaluation protocol and simulation experiments to assess the MHC’s performance in executing diverse behaviors from full and partially-specified target motions.
-We compare the MHC with baseline approaches and analyze the impact of training dataset diversity as well as learning curriculum.
-Additionally, we discuss qualitative observations and the successful sim-to-real transfer of the MHC through real-world trials on the Digit V3 humanoid robot.
-Motion Directive Dataset.
-We used IK retargetting to generate 75 kinematic motion trajectories selected from our motion sources as described in Section 4.
-We aimed to select a diverse set of motions, while avoiding aggressive motions that involve highly dynamic behavior such as jumping, kicking, and large sudden swinging of limbs.
-In order to test models on motions outside of their training sets, we divided the data into three sets, each containing a mix of motions from the original motion sources.
+Experimental Results
+We present our evaluation protocol and simulation experiments to assess the MHC’s performance.
+We compare the MHC with baseline approaches, analyze the impact of varying training sets, and learning curriculum.
+Additionally, we present qualitative and quantitative results of sim-to-real transfer to the Digit V3 robot.
+Component Range
+x velocity ($v_{ix}$) [-0.5, 2.0]m/s
+y velocity ($v_{iy}$) [-0.5, 0.5]m/s
+turn rate ($w_{i}$) [-0.5, 0.5]radians/s
+torso x orientation ($b_{ix}$) [-0.314, 0.314]radians
+torso y orientation ($b_{iy}$) [-0.314, 0.314]radians
+torso height ($b_{ih}$) [0.5, 1.0]m
+Standing mask bit ($I_{is}$) {0, 1}
+Table 1: Predefined ranges for components of locomotion directives.
+Velocities ($v_{ix},v_{iy}$) and turn rate ($w_{i}$) are set to zero when the standing mask bit ($I_{is}$) is set to 11, indicating a standing directive.
+Motion Directive Dataset
+We used IK retargetting to generate 75 diverse kinematic motion trajectories.
+We selected this set of motions based on considering diversity, fit to the robot morphology, and avoiding overly aggressive motions such as jumping, kicking, and high velocity swinging of limbs.
+In order to study the generalization capabilities of our approach, we test MHC models on motions outside of their training sets, dividing the data into three sets, each containing a mix of motions from the original motion sources.
   • setA [20 motions]: Amass Boxing (5), Amass misc (9), Reallusion (2), optimized (2), video (2)
@@
   • setC [35 motions]: Amass Boxing (19), Amass misc (12), Reallusion (1), optimized (3)
-While each set contains some basic motions related to hand waving and simple upper body movements, the general trend is an increasing level of difficulty as judged by the authors.
-We note that setC has a number of the most difficult motions, involving highly dynamic motion, such as tennis smashes and difficult boxing moves.
-The feasibility of these more difficult motions is unclear for our realistic Digit model.
-Experiment Setup and Metrics.
-Each experiment involves evaluating an MHC controller on one of the retargeted motion datasets.
-For each motion we create 3 motion directives corresponding to a whole body directive, upper body command with lower body standing directive, upper body command with lower body locomotion directive.
-For each motion directive we evaluate the MHC over 200 episodes, each having a maximum time of 50 seconds (2500 steps).
-Each episode starts from a randomized point in the directive and then repeatedly cycles through the directive three times before resetting to another randomly selected starting point.
-Based on these episode runs we evauate the following metrics, which depend on the type of directive.
-  • Failure Rate (Fail %): Percentage of episodes that result in failure (i.e. falling) before 50 seconds.
-  • Mean per joint positional error ($\mathbb{E}_{\text{MPJPE}}$): Mean positional error for end effectors (hands, elbows, knees, feet) is calculated as the L2 norm of the current Euler position relative to the torso, compared to values from the directive.
+While each set contains simple upper-body movements (such as waving), the general trend is an increasing level of difficulty as judged by the authors.
+We note that setC has a number of the most difficult motions, involving highly dynamic motion, such as tennis smashes and difficult boxing moves requiring precise and dynamic footwork combined with upper body motions.
+The feasibility of these more difficult motions is not obvious for our realistic Digit model.
+Experimental Setup and Metrics
+We evaluate the MHC on different input modalities including partial/whole-body mimicry, walking, and standing with perturbations.
+We use the full re-targeted dataset for whole-body mimicry and mask out the lower body targets for partial-body mimicry.
+Each episode for full/partial mimicry experiments command the complete motion directive once.
+For pure walking experiments the directives are sampled from a predefined range as specified in Table 1.
+For each walking experiment episodes last for 10s and all non-locomotion directive components are masked out.
+Finally, for standing perturbation testing we follow prior work (van Marum et al.
+2024) and provide an input directive of standing and apply varying forces at the base of the torso for a window of length of 25 steps to measure the root position drift.
+Below we describe the metrics used throughout our evaluations.
+  • Mean End-Effector Positional Error ($\mathbb{E}_{\text{MEPE}}$): Mean positional error for end effectors (hands, elbows, knees, feet) is calculated as the L2 norm of the current Euler position relative to the torso, compared to values from the directive.
 For partial directives, the error is computed only for unmasked joints.
-This is done for both partial and fully-specified directives before episode termination.
   • Root Drift (RootΔ): The mean of drift in root position during an episode from its commanded position when following partially or fully-specified directives.
-Computed as an L2 norm of current root position against expected root position in previous directive step.
-Model Fail % $\mathbb{E}_{\text{MPJPE}}$ RootΔ
-Locomotion+Offsets 49.213 0.156 0.168
-LocomotionTrain 4.800 0.227 0.203
-MHC 0.450 0.088 0.134
+Computed as an L2 norm of the current root position against expected root position in the previous directive step.
+  • Failure Rate (Fail %): Percentage of episodes that result in failure (i.e. falling).
+Comparison to Baselines
+Since no prior approaches handle masked directive inputs like the MHC, we compare our model with different baselines for locomotion and mimicry.
+For locomotion, we compare against a state-of-the-art standing and walking controller(van Marum et al.
+2024).
+We also develop two natural baselines to compare against for locomotion + upper-body mimicry: 1)OffsetBlind.
+This model focuses on solely tracking locomotion components of input directives and does not receive upper-body IK offsets as model input.
+Rather it feeds upper-body offsets from the directive straight to the PD controllers without considering them in its locomotion planning.
+2) OffsetAware.
+This model is the same as OffsetBlind, except that it receives upper-body offsets in its input.
+This allows it to adjust its locomotion control based on the anticipated upper body motions.
+Model Walk RootΔ (m) Stand RootΔ (m)
+1x m/s 0.5y m/s 50N 100N
+MHC 0.413 0.343 0.256 0.311
+SaW(SOTA) 0.803 0.388 0.109 0.452
+Table 2: Locomotion performance comparison of MHC models with the SOTA Stand-and-Walk (SaW) controller.
+Root drift ($\mathrm{Root_{\Delta}}$) is measured for 10 s walking episodes and standing under applied torso forces.
+Model Fail % $\mathbb{E}_{\text{MEPE}}$(in m) RootΔ(in m)
+OffsetBlind 49.213 0.156 0.168
+OffsetAware 4.800 0.227 0.203
+MHC 0.450 0.098 0.141
+Real World 5.0 0.117 -
 […]
-Comparison to Baselines.
-There are no existing approach that handle the masked directive inputs of the MHC.
-Thus, we have developed two baselines to compare against: 1) Locomotion+Offset.
-This controller is trained on the first two curriculum stages and only sees the locomotion commands at its input.
-At inference time, it receives locomotion commands from the directives and for other unmasked joints directly uses the IK offsets in the target motions as the PD setpoints, 2) LocomotionTrain.
-This is the same as Locomotion+Offset during training and inference, but includes the IK offsets as part of the controller input.
-The motivation for LocomotionTrain is that by seeing the IK offsets that will be used during motion generation, the controller can adjust to account for the upcoming motion and be more robust.
-We train the MHC and two baselines on the SetA data.
-The evaluation results are provided in Table 1.
-We see that the MHC outperforms the baselines across all metrics.
+For this comparison experiment we train the MHC and all baselines on datset SetA.
+The results of our standing and walking evaluation in comparsions to the SOTA (van Marum et al.
+2024) are provided in Table 2.
+We give results for the root drift when given walking commands (+1m/s x and +0.5m/s y) and for standing under different force perturbations (50N and 100N).
+The MHC demonstrates competitive performance in both walking and standing tasks, highlighting that MHC is on par with SOTA walking and standing capabilities.
+Table 3 gives results comparing the MHC to the mimicry baselines.
+The MHC outperforms the baselines across all metrics.
 Most significantly it achieves a failure rate of less than 1% which is significantly lower than the baselines.
-Interestingly, we see that LocomotionTrain achieves a significantly lower rate than Locomotion+Offsets, showing the utility having the controller learn to prepare for upcoming target motions.
+Interestingly, we see that OffsetAware achieves a significantly lower failure rate than OffsetBlind, showing the utility of explicitly training the the controller to prepare for upcoming target motions provided in its input.
+Tables 2 and 3 evaluate complementary aspects of MHC performance.
+Table 2 isolates locomotion robustness, comparing MHC against the state-of-the-art SaW controller under force perturbations.
+Table 3 focuses on whole-body imitation with locomotion integration, comparing MHC to mimicry baselines.
+Evaluation of Generalization Performance
 […]
@@
 […]
-Table 2: Comparison of Models across Different Datasets.
-Note that higher failure rates tend to be biased toward smaller (better) values of the other metrics, since those metrics are only averaged over executions before failure.
-Figure 2: Comparison of curriculum
-stage 1 vs. training from scratch
-Figure 3: Comparison of MHC vs. training whole-body directives without stability stage
-Overall, these results show that our training approach can generalize to motions outside of the training set.
-However, more investigation is needed to understand the significance of training and testing with more difficult motions.
-In particular, the exploration of much longer training runs and continually monitoring of individual motion performance is likely to provide useful insights.
-Curriculum Results.
-To evaluate the effectiveness of our curriculum design, we conducted an ablation study comparing the MHC trained with our complete curriculum against two variants: (1) NoCurriculum, a controller trained without any curriculum, and (2) LocoMask, a controller trained without the stabilization stage.
-All controllers were trained on setA.
-The results, shown in Figure 2 and Figure 3, demonstrate the importance of our structured curriculum.
-The NoCurriculum controller took significantly longer to learn and did not reach the same performance level as the MHC in the initial locomotion stage (Figure 2).
-The LocoMask controller exhibited greater instability, resulting in shorter average episode lengths (Figure 3).
-These findings highlight that our staged curriculum, which progressively learns locomotion, stability, and whole-body motion tracking, is very important for achieving our observed robustness and stability.
-Sim-to-Real Results.
-We conduct a real-world demonstration of the MHC using the Digit V3 humanoid robot.
-Our supplementary video illustrates a number of successful examples and some failures.
-Some examples of successes include tracking of fully specified boxing directives, partial directives such as robust walking using joystick commands, box pick and place with locomotion + upper body partial directives as well as non-traditional zombie walking gaits.
-Failure modes included dancing motions that involved balancing on one feet for a substantial amount of time and fully extended arm poses for box pick up motions.
-Here holding the box outwards, away from the torso caused imbalance leading to a continuous drift from standing posture in order to stabilize.
-Wheile there is significant room to improvement these sim-to-real results demonstrate that the MHC framework is a promising path toward highly robust and versatile whole-body humanoid control.
-Figure 4: A) Locomotion Directives to follow joystick commands; B) Fully-specified directive for a boxing jab displaying whole body torso motion and feet coordination; C) Masked partial directives for upper body motion tracking and lower body locomotion commands to achieve the task of moving to a location, picking up the box, moving while holding the box and placing it at a location.
-6 Limitations
-One of the major current limitations is a significant sim-to-real gap.
-This is particularly prevalent for motions where the feet are placed widely or where one foot remains in the air for an extended period.
-Our hypothesis is that this sim-to-real limitation requires research that is more generally focused on sim-to-real transfer, possibly via the incorporation of real-world data.
-The MHC has been a useful vehicle for highlighting these gaps.
-Another limitation of our current training approach is that we only consider a limited set of masking patterns during training.
-Supporting fully general masking may lead to additional robustness and open up new possibilities for authoring behaviors.
-Finally, the MHC is currently does not account for interaction with external objects, such as the box, during a box pickup.
-Explicitly incorporating objects to be manipulated into the MHC input and training is an important direction for supporting practical applications.
-Appendix A Dynamics Randomization
-All models in our work were trained with a fixed dynamics randomization setup in order to transfer from simulation to real without learning the exact dynamics of the system.
-The ranges for all elements are given in Table 3.
-Element Range(%)
-Joint Damping [-50, 250]
-Body Mass [-25, 25]
-CoM Position [-5, 5]
-Spring Stiffness [-10, 10]
-Friction [-20, 20]
-Range(radians)
-Encoder Noise [-0.05, 0.05]
-Ground Slope [-0.05, 0.05]
-Table 3: Dynamics Randomization Ranges.
-Appendix B Implementation Details
-MHC has been trained using actor critic based PPO algorithm. Table 4 specifies the hyperparameters used for each training cycle independent of the curriculum.
-Hyperparameter Value
-Learning Rate (Actor & Critic) 3e-4
-Buffer Size 50000
-Number of Epochs 5
-Batch Size (of episodes) 8
-PPO Log-Clip Range 0.2
-Maximum Gradient Norm 0.05
-Discount Factor (Gamma) 0.95
-GAE Lambda 0.95
-Table 4: PPO Hyperparameters
-Appendix C Reward structure
-Table 5 gives a detailed view of our complete reward function. Note that style rewards are only ignored when following a fully-specified directive denoted by $I_{i}==1$.
-Following notation has been used:
-  • $i$ indicates current time step
-  • $I_{i}==1$ indicated if all joints are being mimicked and its a whole body motion directive;
-  • $v_{xy}$ = base linear velocity;
-  • $s$ = current robot parameter;
-  • $\boldsymbol{b}$ = base acceleration;
-  • $c_{feet}$ = default foot position xyz and orientation rpy;
-  • $qd(\cdot)$ = quaternion distance function;
-  • $n_{c}$ = number of feet in contact with ground;
-  • $\mathds{1}_{td}$ = boolean variable indicating a foot touchdown in the current timestep;
-Reward Term Definition Weight Active if $I_{i}==1$
-if $v_{i}==0$ [Standing] else [Walking]
-Task $x,y$ velocity $\exp\left(-5\cdot(v_{xy})\right)$ $\exp\left(-5\cdot(v_{xy}-v_{i})^{2}\right)$ 0.15
-Tracking Reward $\exp\left(-1.5\cdot||(\theta-\hat{\theta}_{i})\cdot I_{i}||\right)$ 0.2
-Task, Style Yaw orient $\exp\left(-300\cdot qd(\omega_{z},w_{i})\right)$ 0.1
-Roll, pitch orient $\exp\left(-30\cdot qd\big{(}\omega_{xy},(b_{i}[xy])\big{)}\right)$ 0.2
-Base height $\exp\left(-20\cdot\big{|}s_{z}-(b_{i}[z])\big{|}\right)$ 0.05
-Style Feet orientation $\exp\left(-\sum|s_{\text{feet, rpy}}-c_{\text{feet, rpy}}|\right)$ 0.05
-Feet position $\exp\left(-3\cdot|S_{\text{feet, xyz}}-c_{\text{feet, xyz}}|\right)$ 1 0.05
-Feet airtime 1 $\sum_{f\in(l,r)}(t_{air,f}-0.4)*\mathds{1}_{td,f}$ 1.0
-Feet contact 1 $\begin{cases}1&\text{if }n_{c,t^{*}}=1\text{ for any }t^{*}\in[t-0.2,t]\\ 0&\text{else}\end{cases}$ 0.1
-Arm $\exp\left(-3\cdot\big{|}\big{|}(\theta[\hat{J}]\text{ where }\hat{J}\in\text{ arms})\big{|}\big{|}\right)$ 0.03
-Reg Base acceleration $\exp\left(-0.01\cdot\sum|b_{xyz}|\right)$ 0.1
-Action difference $\exp\left(-0.02\cdot\sum|a_{t}-a_{t-1}|\right)$ 0.02
-Torque $\exp\left(-0.02\cdot\frac{1}{N}\sum|t_{\text{motor}}|/t_{\text{max}}\right)$ 0.02
-Table 5: Reward Terms
+Metric $\mathbb{E}_{\text{MEPE}}$(in m) RootΔ(in m)
+setA setB setC setA setB setC
+[…]
+[…]
+[…]
+Table 4: Performance Comparison of MHC models on partial and whole-body directives computed on different datasets (setA, setB, setC)
+Curriculum Results
+We also train and evaluate $\text{MHC}_{\text{NC}}$ that is trained without any curriculum on the full set of directives (locomotion, partial and whole-body).
+We trained $\text{MHC}_{\text{NC}}$ until the training loss plateaued for a long period of time, providing substantially more training time than the MHC.
+In Table 5 we compare $\text{MHC}_{\text{NC}}$’s performance against MHC in perturbed standing, walking, and mimicry.
+We see that MHC clearly outperforms $\text{MHC}_{\text{NC}}$ on the tasks of walking and mimicry, which shows the effectiveness of our curriculum design.
+It is interesting to note that $\text{MHC}_{\text{NC}}$ slightly outperforms the MHC with respect to standing against perturbation.
+This appears to come with the trade-off of not aligning itself to the command directive as well as the MHC, particularly in the lower body.
+Figure 2: Real-world demonstrations of our approach.
+A) Locomotion Directives using joystick commands.
+B) Fully-specified directive for a boxing jab; showing whole-body torso motion and foot coordination.
+C) Handcrafted sequence of masked directives combining upper body motion trajectory and lower body joystick commands; showing the ability to move to a location, pick up a box, move while holding it, and place it at a target location.
+Metric MHC $\text{MHC}_{\text{NC}}$
+Stand (50 N) RootΔ(in m) 0.256 0.189
+(100 N) RootΔ(in m) 0.311 0.234
+Walk (1x m/s) RootΔ(in m) 0.413 1.346
+(0.5y m/s) RootΔ(in m) 0.343 0.937
+$\mathbb{E}_{\text{MEPE}}$(in m) 0.098 0.149
+Mimic RootΔ(in m) 0.141 0.327
+Fail % 0.450 4.875
+Table 5: Performance metrics comparing a model trained with curriculum (MHC) against a model trained without curriculum ($\text{MHC}_{\text{NC}}$) on the tasks of standing, walking and mimicry (averaged for both partial and whole-body directives from motions in dataset setA).
+Sim-to-Real Results
+We conducted real-world demonstrations of the MHC using the Digit V3 humanoid robot.
+Figure 2 shows trials including box locomanipulation, which involved the input modality of locomotion and upper-body end-effector tracking.
+For a quantitative evaluation we tested the 20 motions in setA on the real robot, performing 3 repetitions of each motion.
+The results are given in Table 3.
+While the failure rate (5%) and joint position error are higher than observed in simulation, these results indicate our approach is able to achieve highly non-trivial sim-to-real transfer.
+Future Work
+This work motivates a number of important research directions: 1) Addressing the remaining sim-to-real gaps, especially for motions with wide foot placements or extended single-foot balancing stances.
+Incorporating real-world data into the sim-to-real process is likely to play an important role.
+2) Continuing to increase the number and types of motions used for MHC training.
+Currently, for Digit V3, the bottleneck is our reliance on the MuJoCo simulator, compared to other higher throughput simulators such as Isaac Sim, which do not yet support the Digit V3 model due to modeling complexities.
+3) Increasing the variety of masking patterns used for training to further enhance versatility.
+4) Handling inconsistent and/or impossible input directives (e.g.
+root velocity is inconsistent with other joint components), which are currently not part of training.
+Acknowledgments
+This work is supported by NSF Award 2321851 and DARPA contract HR0011-24-9-0423.